ln of 5 - lnx equals 3
Answer by MathLover1(20855)


ln(5) + ln(x)=3
ln(5 * x)=3


e^(ln(5x)) = e^3
e cancels ln so..
5x = e^3...................e is an irrational constant approximately equal to 2.718281828.
5x = (2.718281828)^3.
5x =20.09
x =20.09/5
x = 4.018